owenjiang:

use absolute value to express the distance between -10 and 16 on the number line.

dontsaymyname:

What are the answer choices?

Shadow:

\[|-10| + 16 = ?\]

jimthompson5910:

Distance from A to B = |A - B| or Distance from A to B = |B - A| For instance, Distance from -12 to 7 = |-12-7| = |-19| = 19 Distance from -12 to 7 = |7-(-12)| = |19| = 19 The answer to your question isn't 19 as this is just an example problem. As an alternative, you can count out the number of spaces between the two numbers on the number line.